OpenStudy (anonymous):

Evaluate the following indefinite integrals: cos x sin 5x dx

OpenStudy (anonymous):

Angle sum/difference identity for sine: \[(1)~~\sin(x+y)=\sin x\cos y+\sin y\cos x\\ (2)~~\sin(x-y)=\sin x\cos y-\sin y\cos x\] Observe that adding the two together gives the following: \[\sin(x+y)+\sin(x-y)=2\sin x\cos y\] \[\frac{\sin(x+y)+\sin(x-y)}{2}=\sin x\cos y\] So for this integral, rewrite by replacing \(x\) with \(5x\) and \(y\) with \(x\): \[\int\cos x\sin5x~dx=\frac{1}{2}\int\left(\sin6x+\sin4x\right)~dx\]